OpenStudy (anonymous):

A tool box has a volume of x3 + 8x2 + 11x – 20 cm3 and the height is x + 5 cm. Find the polynomial that would represent the area of the bottom of the tool box? Explain your reasoning.

OpenStudy (anonymous):

your real job is to divide \[\frac{x^3+8x^2+11x-20}{x+5}\] i would use synthetic division,but you can do it with long division

OpenStudy (anonymous):

you will get \[x^2+3x-4\] do you need the steps?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

do you know how to use synthetic division?

OpenStudy (anonymous):

synthetic division may be easier, but i think it's bad for the environment :)

OpenStudy (anonymous):

it is a better environment for synthetic than long division. i can write the process but if you don't know it you will not understand it

OpenStudy (anonymous):

No, can you show me how you would do it?

OpenStudy (anonymous):

list the coefficients of \[x^3+8x^2+11x-20\] they are 1 8 11 -20

OpenStudy (anonymous):

because you are dividing by x + 5 put -5 on the side 1 8 11 -20 -5 _______________________________

OpenStudy (anonymous):

bring down the 1 1 8 11 -20 -5 _______________________________ 1

OpenStudy (anonymous):

1*-5=-5 1 8 11 -20 -5 -5 _______________________________ 1

OpenStudy (anonymous):

8 - 5 = 3 1 8 11 -20 -5 -5 _______________________________ 1 3

OpenStudy (anonymous):

3*-5=-15 1 8 11 -20 -5 -5 -15 _______________________________ 1 3

OpenStudy (anonymous):

11 - 15 = -4 1 8 11 -20 -5 -5 -15 _______________________________ 1 3 -4

OpenStudy (anonymous):

-5 * -4 = 20 1 8 11 -20 -5 -5 -15 20 _______________________________ 1 3 -4

OpenStudy (anonymous):

-20 + 20 = 0 1 8 11 -20 -5 -5 -15 20 _______________________________ 1 3 -4 0

OpenStudy (anonymous):

your solution is the bottom row. it is \[x^2+3x-4\]

OpenStudy (anonymous):

so what hapened to the last part? explain ur reasoning?